1. What is the total return to an investor who purchases a bond for $1000 and sells the bond for $1,041 next year. Assume the bond has an annual coupon rate of 4% that is paid in two equal payments.

Record your answer as a decimal to four places after the decimal, so if your answer is 4.212111%, record your answer as 0.0421.

2. A 9-year zero coupon bond has a yield to maturity of

5.5 percent, and a par value of $1,000.  What is the price of the bond?

3. A 12-year bond has a 9 percent annual coupon, a yield to maturity of
4.2 percent, and a face value of $1,000.  What is the price of the bond?

4. Interest rate risk is associated with the bonds price variability given a change in the interest rates.

Suppose you have BOND A, which is a 30 year zero coupon bond and BOND B, which is a 5 year 10% coupon bond. If interest rates (YTM) change from 8% to 7% the bonds will increase in value. Suppose BOND A's price changes from $99.38 to 117.56 and the 5 year 10% coupon bond price changes from $1079.85 to $1123.01. Which bond has the greatest percentage increase in value? Record the percentage increase in value of the bond with the highest percentage change below. Write the increase as a decimal, so a 5% increase would be written as 0.0500.

PLEASE ANSWER ALL QUESTIONS
Answer #1

1.
=(1041+4%*1000)/1000-1=8.100%

2.
=1000/1.055^9=617.629261296125

3.
=90/4.2%*(1-1/1.042^12)+1000/1.042^12=1445.300756117

4.
% change in Bond A=117.56/99.38-1=18.293%